For the reaction shown, calculate how many moles of each product form when the given amount of each reactant completely reacts. Assume that there is more than enough of the other reactant. C3H8(g)+5O2(g)→3CO2(g)+4H2O(g)

Express your answer using two significant figures.

A- 14.7 molC3H8

B- 4.7 molC3H8

C- 0.0595 molC3H8

D- 4.7 molO2
Answer #1

1 mole of C3H8 gives 3 moles of CO2 and 4 moles of H2O (as per given stoichiometric equation)

A. 14.7 mole C3H8

moles of CO2 = 14.7 * 3= 44.1 moles of CO2

moles of H2O = 14.7 * 4= 58.8 moles of H2O

B. 4.7 mole C3H8

moles of CO2 = 4.7 * 3= 14.1 mole of CO2

moles of H2O = 4.7 * 4 = 18.8 moles of H2O

C. 0.0595 mole C3H8

moles of CO2 = 0.0595 * 3 = 0.1785 mol of CO2

moles of H2O = 0.0595 * 4 = 0.238 moles of H2O

D. 4.7 mol O2

(5 moles O2 gives 3 moles CO2 and 4 moles H2O

That means 1 mole O2 gives 3/5= 0.6 moles of CO2 and 4/5= 0.8 moles of H2O)

moles of CO2 = 4.7 * 0.6 = 2.82 moles of CO2

moles of H2O = 4.7 * 0.8= 3.76 moles of H2O